OpenStudy (anonymous):

integrate dx/(4+x^2)^2 using trig substituition. im confused i thought it had to be a square root to be used no squared can some explain how toset these up?

zepdrix (zepdrix):

\[\large\rm \int\limits \frac{dx}{(4+x^2)^2}\]Ummm trig sub? Yah that'll be fun :)

zepdrix (zepdrix):

If you factor a 4 out of each term in the denominator you get\[\large\rm \int\limits\limits \frac{dx}{(4+x^2)^2}=\frac{1}{4^2}\int\limits\frac{dx}{\left(1+\frac{1}{4}x^2\right)}\]Let's bring the 1/4 into the square with the x,\[\large\rm =\frac{1}{16}\int\limits\limits\frac{dx}{\left(1+\left[\frac{x}{2}\right]^2\right)}\]Now we have that thing in the form: \(\large\rm 1+stuff^2\) So we can make the substitution: \(\large\rm stuff=\tan\theta\)

zepdrix (zepdrix):

\[\large\rm \frac{x}{2}=\tan\theta\]From there, find your \(\large\rm d\theta\). Plug in all the stuff. Then we'll use some cool trig rules to simplify everything down, and hopefully, be able to integrate easily.

zepdrix (zepdrix):

From there, find your \(\large\rm dx\)* is what I meant to say. Any trouble finding it? :o

zepdrix (zepdrix):

Or were my steps confusing?

OpenStudy (anonymous):

Im sorry I just did not understand the find dtheta part. I converted (1+tan0) to sec^20 can i take the derivative of that to dnd d0?

zepdrix (zepdrix):

\[\large\rm =\frac{1}{16}\int\limits\limits\limits\limits\frac{dx}{\left(1+\left[\color{orangered}{\frac{x}{2}}\right]^2\right)^{2}}=\frac{1}{16}\int\limits\limits\limits\limits\frac{dx}{\left(1+\color{orangered}{\tan^2\theta}\right)^{2}}=\frac{1}{16}\int\limits\frac{dx}{\sec^4\theta}\]So you've got that much figured out so far? To find dx, you need to go back to your original substitution. \(\large\rm \frac{x}{2}=\tan\theta\) Take derivative of that thing.

OpenStudy (anonymous):

so it would turn do x=2tantheta and the derivative is sec^2theta do it would be dtheta over sec^2theta

zepdrix (zepdrix):

Mmm are you missing a 2 in your dx maybe? Should look like this.\[\large\rm x=2\tan theat\qquad\to\qquad \color{royalblue}{dx=2\sec^2\theta~d\theta}\] \[\large\rm \frac{1}{16}\int\limits\limits\frac{\color{royalblue}{dx}}{\sec^4\theta}=\frac{1}{16}\int\limits\limits\frac{\color{royalblue}{2\sec^2\theta~d\theta}}{\sec^4\theta}\]

zepdrix (zepdrix):

Ah I made a typo in my x :) lol

zepdrix (zepdrix):

\[\large\rm =\frac{2}{16}\int\limits\frac{d \theta}{\sec^2\theta}\]So something like this? Ya you have the right idea.

zepdrix (zepdrix):

Recall that \[\large\rm \cos\theta=\frac{1}{\sec\theta}\]

OpenStudy (anonymous):

yeah i have that! and i would just u sub right?!

zepdrix (zepdrix):

\[\large\rm =\frac{1}{8}\int\limits \cos^2\theta~d \theta\] Hmm no. Looks like you'll have to use your `Half-Angle Formula` at this point. :)

OpenStudy (anonymous):

i got 1/8 * integral of cos^2theta

OpenStudy (anonymous):

\[=\frac{ 1 }{ 16 }\int\limits \left( 2\cos ^2 \theta \right)d \theta=\frac{ 1 }{ 16 }\int\limits \left( 1+\cos 2\theta \right) d \theta =?\]

OpenStudy (anonymous):

its going to be theta -sin2theta/2 but that would make the denominator for sin 32 and it is wrong @surhithayer

zepdrix (zepdrix):

No, that sounds fine :) You just need your final answer in terms of \(\large\rm x\), not \(\large\rm \theta\). So you need to undo your substitution, perhaps with the use of a triangle.

OpenStudy (anonymous):

\[\int\limits \cos 2\theta d \theta =\frac{ \sin 2\theta }{ 2}\]

OpenStudy (anonymous):

\[=\frac{ 1 }{ 2 }*\frac{ 2 \tan \theta }{ 1-\tan ^2 \theta }=\frac{ \tan \theta }{ 1-\tan ^2 \theta }\]